OpenStudy (anonymous):

what is the slope of -x+5y=4 a.1/5 b.-5

Mehek (mehek14):

first, move x to the other side

Mehek (mehek14):

you have to turn it into the general formula \(\Large{y = mx + b}\)

Mehek (mehek14):

m = slope

Mehek (mehek14):

can you do that? @rhondahyde16

OpenStudy (anonymous):

i dont understand how to do it with the equation i got

Mehek (mehek14):

you have -x + 5y = 4 move -x to the other side

OpenStudy (anonymous):

so " 5y=4+x"..? :/ ( stuff like this really confuses me)

Mehek (mehek14):

yup

Mehek (mehek14):

now we need to get rid of 5 from y

Mehek (mehek14):

so divide 5 with 5y, x, and 4

OpenStudy (anonymous):

i got 0.8, wrong? or no.

Mehek (mehek14):

hm 5y/5 = y x/5 4/5 so y = x/5 + 4/5

Mehek (mehek14):

x/5 is equivalent to 1/5x

Mehek (mehek14):

so y = 1/5x + 4/5

Mehek (mehek14):

what's the slope?